### v1

# CogGuide: Human-Like Guidance for Zero-Shot Omni-Modal Reasoning

Abstract—Targeting the issues of “shortcuts” and insufficient contextual understanding in complex cross-modal reasoning of multimodal large models, this paper proposes a zero-shot multimodal reasoning component guided by human-like cognitive strategies centered on an “intent sketch” The component comprises a plug-and-play three-module pipeline—Intent Perceiver, Strategy Generator, and Strategy Selector—that explicitly constructs a “understand–plan–select” cognitive process. By generating and filtering “intent sketch” strategies to guide the final reasoning, it requires no parameter fine-tuning and achieves cross-model transfer solely through in-context engineering. Information-theoretic analysis shows that this process can reduce conditional entropy and improve information utilization efficiency, thereby suppressing unintended shortcut reasoning. Experiments on IntentBench, WorldSense, and Daily-Omni validate the method’s generality and robust gains; compared with their respective baselines, the complete “three-module” scheme yields consistent improvements across different reasoning engines and pipeline combinations, with gains up to approximately +9.51 percentage points, demonstrating the practical value and portability of the “intent sketch” reasoning component in zero-shot scenarios.
Index Terms—LLM, Omni-Modal, Intent Sketch, Information Entropy
Figure 1. Relative Percentage Improvements over Baselines by the Proposed Method on IntentBench, WorldSense, and Daily-Omni
INTRODUCTION
In recent years, multimodal large language models (MLLMs) have been propelling artificial intelligence from purely textual understanding toward comprehensive understanding and reasoning across “all modalities,” including vision, speech, and video. In 2023, OpenAI released GPT-4, which for the first time introduced visual input capabilities, enabling LLMs to parse images and perform reasoning [1]. Since then, numerous researchers have begun developing and applying multimodal large models [2], [3]. Against this backdrop, tasks such as video question answering impose higher demands on dynamic-scene comprehension and temporal reasoning[4], motivating research to unify, align, and fuse vision, language, and audio in order to enhance cross-modal understanding and reasoning [5]. Consequently, multimodal large models have expanded from images to video and audio, evolving toward a general “audiovisual omnipotent” architecture that strengthens world representations while broadening human–computer interaction [6].
However, existing multimodal models still have shortcomings in complex reasoning scenarios. Even with massive parameters and multimodal pre-training, they continue to face challenges such as insufficient global contextual understanding and “shortcut” reasoning in complex tasks [7]. Models often over-rely on local or single-modality cues while overlooking critical cross-modal information, leading to outputs that deviate from human intent [8]. On the other hand, even when employing prompting strategies such as Chain-of-Thought, current multimodal large models remain notably limited on multi-step cross-modal reasoning tasks [9]. Constructing reasoning chains via reinforcement learning may cause models to acquire “shortcut” strategies, thereby reducing generalization [10]. These phenomena indicate that relying solely on the model’s inherent reasoning ability and simple prompts is still insufficient to align the model with human intent.
To address these issues, researchers have begun incorporating “intent” as a mediating variable connecting user queries and cross-modal evidence into the multimodal reasoning loop: one line of work introduces explicit intent labels or scene purposes in multimodal question answering and video understanding to constrain candidate answers and the scope of reasoning [7]; another employs instruction tuning and templated prompts to explicitly declare desired behaviors on the input side, thereby implicitly aligning reasoning goals [11]; still others construct “intent-conditioned” retrieval–reasoning pipelines or agents so that evidence selection and reasoning steps are driven by the current intent [12] ; there are also methods that adopt text-guided multimodal fusion to assist multimodal intent understanding [13]. Meanwhile, attempts to exploit the temporal and event structures of audio–video to infer latent intent and filter irrelevant cues are increasing [14], [15]. However, these methods have certain limitations: they rely on dense annotation and task-specific training, making zero-shot transfer difficult [7], [16], or they treat intent as static labels or prompt fragments without externalizing it into generable, assessable, and selectable strategies, thereby failing to stably suppress “shortcut” reasoning and local biases [8], [11].
Recently, studies have begun to explore multimodal reasoning that simulates human “sketches of thought.” For example, the Sketch-of-Thought framework [17] introduces a human-like cognitive reasoning paradigm that maintains reasoning accuracy while reducing verbose intermediate reasoning; the Machine Mental Imagery approach [18] enables models to incorporate latent visual “imagination” representations during reasoning, simulating the human process of using mental imagery to aid reasoning, thereby improving reasoning efficiency and understanding of complex scenes. Inspired by such research, we propose a human-like cognitive intent-sketch reasoning component that modularly enhances multimodal intent recognition and reasoning performance. The core idea is to decompose and recombine the context of complex problems and to embed human cognitive processes into model reasoning through explicit strategy-planning prompts (i.e., the “intent sketch” mechanism), thereby strengthening the model’s understanding of multimodal information and latent intent. Specifically, we design a plug-and-play three-module pipeline: Intent Perceiver, Strategy Generator, and Strategy Selector. The Intent Perceiver module analyzes video scenes and generates textual intent; the Strategy Generator module produces multiple candidate reasoning strategies based on the context; and the Strategy Selector module evaluates these candidate strategies and selects the optimal one. Finally, a large language model (serving as the reasoning engine) combines the selected strategy with the context and produces an answer by injecting prompts. This component fully leverages the idea of in-context engineering: the entire process requires no parameter fine-tuning of the underlying models; instead, it dynamically injects high-quality strategy prompts to guide the model to focus on the correct lines of thought and cues during reasoning. Under a zero-shot setting, our method can rapidly adapt to different pretrained multimodal models, achieving general performance gains across models. On multiple challenging multimodal reasoning benchmarks (e.g., IntentBench [7], WorldSense[14] and Daily-Omni[15]), zero-shot evaluations have been conducted by prior work [7]. Taking Qwen2.5-Omni as an example, its baseline accuracies on IntentBench, WorldSense, and Daily-Omni are 64.2%, 45.4%, and 47.45%, respectively, serving as a unified comparison baseline. Our method surpasses the original model on all benchmarks (with accuracy improvements of up to 9.51%). Meanwhile, in terms of “effectiveness vs. cost,” compared with training-based methods and without parameter updates or data preparation, we maintain zero training overhead and low latency, resulting in better overall cost and deployment timelines.
The main contributions of this paper are as follows:
  • Human-like cognitive intent-sketch reasoning component: We design a cognitive reasoning component comprising three modules (Intent Perceiver, Strategy Generator, Strategy Selector) that simulates the human cognitive process of “understanding context–planning strategies–selecting solutions.” By explicitly planning and selecting contextual strategies, it enhances reasoning accuracy on complex multimodal tasks. The component is dynamically assembled at inference time, can handle multimodal tasks such as vision and audio, requires no supervision signals or parameter fine-tuning, and effectively improves the model’s understanding of complex scenes and latent intent under zero-shot conditions. Meanwhile, this modular “intent sketch” reasoning component can be embedded as a plug-and-play unit into existing multimodal models; by providing high-quality strategy context, it yields performance gains without fine-tuning, demonstrating cross-model generality and portability.
  • Theoretical analysis from an information-entropy perspective: From an information-theoretic standpoint, we provide a detailed analysis of the working mechanisms of each module, examine how strategy prompts reduce uncertainty in model decisions and improve information utilization efficiency, and present related theoretical derivations, thereby deepening understanding of the principles of context optimization in zero-shot reasoning.
  • Comprehensive multimodal experimental validation: We conduct ablation and comparative experiments on multiple multimodal reasoning benchmarks—including the human-intent understanding task IntentBench and the audio–video collaborative analysis datasets WorldSense and Daily-Omni—to verify that the component can improve the reasoning accuracy of multimodal models without additional training. In addition, by substituting different large reasoning models as the reasoning engine, we validate the component’s generality, highlighting the practicality of strategy prompts in zero-shot reasoning and the value of in-context engineering.
Figure 2. Intent sketch guided zero-shot omni-modal reasoning component: three-module pipeline and information-entropy analysis
METHOD
We propose a human-like cognitive intent-sketch reasoning component comprising three serially connected modules: Intent Perceiver, Strategy Generator, and Strategy Selector. The system receives multimodal input $X=(V,A,Q)$, where $V$ denotes a sequence of video frames, $A$ denotes audio, and $Q$ denotes the user’s textual query. The Intent Perceiver compresses $(V,A)$ into a linguistic representation and extracts intent-related cues pertinent to the question; the Strategy Generator, conditioned on the input $\{Q,\text{Intent}\}$, produces $N$ compact and complementary reasoning strategies; the Strategy Selector, given the question (and options, if any), selects from the $N$ candidates the strategy $S^{*}$ that best fits the current question, which then guides the final chain-of-thought reasoning to generate the answer. The entire process mirrors how humans tackle complex tasks—first understanding intent, then planning strategies, and finally selecting the optimal plan—thereby ensuring a deeper grasp of the question and a clearer reasoning trajectory. Ultimately, the selected strategy is used as explicit guidance fused into the model’s reasoning context, thereby influencing and improving the final answer generation.
2.1 Intent Perceiver: Intent Representation for Strategy Generation
The Intent Perceiver module performs comprehensive analysis and cross-modal evidence fusion for multimodal inputs to distill the intent and key information most closely related to problem solving, thereby clarifying “what the model should plan around.” The module ingests multimodal cues such as video and audio, parses them, identifies the core aspects that the user’s question truly targets, and outputs an intent representation oriented toward downstream reasoning, in order to reduce ambiguity and noise at the question-understanding stage and establish a clear starting point for strategy planning and answer generation.
This module forms an intent representation jointly from the multimodal input and the question: let the input be $X=(V,A,Q)$, where $V$ is video, $A$ is audio, and $Q$ is the question text; the module outputs the intent representation $Z_{IP}=f(X)$. From the perspective of Shannon information theory, $Z_{IP}$ is a task-relevant distillation of $X$, mapping from high-dimensional multimodal perception to a discrete semantic representation. The goal is to provide additional information gain for the subsequent generation of strategy $S$, so that the information on which subsequent reasoning relies is more targeted to the question’s objective. Considering two ways of generating strategies—whether to include the intent representation $Z_{IP}$ (input only $Q$ vs. input $Q,Z_{IP}$)—compared with using only $Q$, adding $Z_{IP}$ is equivalent to providing nonnegative information gain, satisfying:
$MI\left(S;\left(Q,Z_{IP}\right)\right)-MI(S;Q)=MI\left(S;Z_{IP}\mid Q\right)\geq\$0 (1)
From an entropy standpoint, this corresponds to $H(S\mid Q,Z_{IP})\leq H(S\mid Q)$. In practice, we treat $Z_{IP}$ as an information carrier at an intermediate “bottleneck,” embodying the fusion of the question and multimodal cues, so that the statistical relations associated with the strategy become more concentrated, thereby reducing uncertainty in subsequent strategy planning and decision-making.
Overall, the Intent Perceiver helps reduce uncertainty at the problem-understanding stage—that is, it diminishes ambiguity and multiplicity in the question semantics (reducing “conditional entropy”). By focusing on the key information within the multimodal input that is relevant to the answer, the module to some extent removes noise unrelated to the question, increases the association between the extracted cues and the final answer $Y$, and thus lowers the model’s effective uncertainty when facing input $X$. In other words, by clarifying problem intent, the Intent Perceiver module reduces the ambiguity at the starting point of problem solving and, in subsequent reasoning, effectively lowers the initial value of the conditional entropy $H(Y\mid X)$, providing a more determinate starting point for the entire reasoning chain.
2.2 Strategy Generator: Multi-Strategy Generation Based on Semantic Entropy
After obtaining the intent representation $Z_{IP}=f(V,A,Q)$, the Strategy Generator invokes a large language model (LLM) conditioned on $Z_{IP}$ and the question $Q$, configuring it as a “solution-strategy provider” to produce $N$ distinct candidate reasoning strategies $\{S_{1},S_{2},\ldots,S_{N}\}$. These strategies specify only the key lines of reasoning and modes of evidence organization without directly giving the answer, effectively providing comparable “sketches of thought.” This step relies entirely on prompt engineering to activate the pretrained model’s internal knowledge and introduces no additional trainable parameters.
From an information-theoretic perspective, the strategy can be viewed as a random variable $S$. To characterize uncertainty at the “strategy layer,” we define a set of semantic equivalence classes $\mathcal{M}$ over the strategy semantic space, merging strategies that differ in expression but are semantically equivalent into the same class $m\mathcal$$\in$ M}. Given $(Q,Z_{IP})$, the model’s posterior over semantic classes is denoted $p(m\mid Q,Z_{IP})$, from which we define the semantic entropy
[…]
Compared with conditioning on $Q$ alone, explicitly incorporating $Z_{IP}$ amounts to providing additional information such that
$MI(S;Z_{IP}\mid Q)\ =\ H_{sem}(S\mid Q)-H_{sem}(S\mid Q,Z_{IP})\ \geq\ 0$ (3)
Thereby reducing uncertainty at the strategy layer and making the generated strategies more attuned to the video-related context while focusing on the key information needed to solve the current problem.
Generating multiple strategies in one pass can be understood as sampling from the posterior $p(S\mid Q,Z_{IP})$ to cover multiple high–probability-mass reasoning hypotheses. Given the candidate set $\{S_{1},S_{2},\ldots,S_{N}\}$, the downstream step will select the strategy that best fits the current task; even before selection occurs, the presence of multiple candidates already expands the explorable reasoning space and increases the chance of finding a strategy that yields smaller conditional uncertainty $H(Y\mid Q,Z_{IP},S)$. To prevent semantic collapse of candidate strategies onto a single path while maintaining diversity and coverage of the strategy set, we unify single-strategy clarity (low entropy) and set-level coverage (high entropy) within an optimization framework. Specifically, for each candidate $S_{i}$, let its posterior over semantic classes be $p_{i}(m)=p(m\mid Q,Z_{IP};S_{i})$; its single-strategy semantic entropy is then
[…]
where $\alpha,\gamma>0$ are weights, and $Div(\cdot)$ can be defined via pairwise distances in semantic similarity between candidates to encourage complementary strategies with different emphases, such as “evidence-first,” “temporal/causal-first,” or “cross-modal-alignment-first.” Through this process, the Strategy Generator provides a structured, comparable set of candidate lines of thought aligned with $(Q,Z_{IP})$ for subsequent strategy selection and answer generation.
Overall, by explicitly introducing the reasoning-path variable $S$ and sampling multiple possible values for it, the Strategy Generator increases the diversity of candidate strategies and avoids the model’s premature confinement to a single path, enabling a more comprehensive exploration of the solution space. This supplies more comparative information for the subsequent steps and thereby lays the groundwork for reducing $H(Y\mid X)$.
2.3 Strategy Selector: Strategy Selection Based on Minimum Conditional Entropy/Bayesian Risk
[…]
Equivalently, the selection can be viewed as maximizing information gain, from the relationship between conditional entropy and mutual information:
$IG(Y;S_{i}\mid X)=H_{\theta}(Y\mid X)-H_{\theta}(Y\mid X,S_{i})$ (7)
so that:
[…]
Moreover, because the candidates are $N$ complementary strategies generated in the previous stage under $(Q,Z_{IP})$, selection under the above criterion is equivalent to solving over a finite set:
[…]
This mechanism ensures that, when the model ultimately answers the question, it follows a problem-solving approach that is demonstrably more feasible and efficient, thereby increasing the specificity and reliability of reasoning. From a Bayesian chain perspective of the entire pipeline, explicitly conditioning on the “selected strategy” contracts the answer posterior from $p_{\theta}(Y\mid X)$ to $p_{\theta}(Y\mid X,S^{*})$. Combined with the analytical framework that “each additional effective condition reduces uncertainty about the final answer,” the strategy-selection step provides a prior constraint on the reasoning path, further concentrating the posterior and reducing decision ambiguity and empirical error.
2.4 A Unified Information-Theoretic Framework of Uncertainty Reduction
[…]
EXPERIMENTS AND RESULTS
3.1 Experimental Settings
We evaluate the proposed method on multiple multimodal reasoning benchmark datasets, including the human-intent understanding task IntentBench, the audio–video collaborative analysis dataset WorldSense, and the daily-life scenario dataset Daily-Omni. Accuracy (%) is used as the evaluation metric to compare the baseline performance of the original models (without any strategy prompts) against the reasoning performance after introducing different combinations of modules. The method consists of three pluggable modules—Intent Perceiver, Strategy Generator, and Strategy Selector—and the final output is fed into a reasoning engine to produce answers (see Table 1, “Model Overview”).
Each experiment is conducted under a zero-shot setting, i.e., no additional fine-tuning of the models and zero-shot prompts are used. We select three large multimodal models as reasoning engines: HumanOmniV2, Qwen2.5-Omni, and Qwen2.5-VL, which represent different categories of multimodal large models. Our method employs four pretrained language models as pipeline module models (Pipeline): the closed-source GPT-4o and Doubao-Seed-1.6, and the open-source GLM-4.5 and Qwen3.
To quantify the contribution of each front-end module to the final reasoning performance, we conduct ablation studies centered on the three modules—Intent Perceiver, Strategy Generator, and Strategy Selector (see Table 2). CG_Qwen_vl indicates the complete three-module configuration: take Qwen2.5-VL-32B as intent perceiver, then three candidate strategies are produced, and the selector is enabled to choose the best; CG_GLM_vl take GLM-4.5V as the intent perceiver model to test the substitutability of the intent perceiver; Abl_No_Intent disables the intent perceiver module and inputs the question text directly into strategy generation; Abl_Single_Strategy sets the number of strategies to 1 and disables the selector, thereby isolating the gain from “multiple strategies + selection”; BaseLine completely removes the front-end pipeline (no intent perceiver, no strategy generation, no selection), and the reasoning engine answers solely based on the original multimodal input and the question. Each group is evaluated under the Cartesian product of the four Pipeline models (GPT-4o / GLM-4.5 / Doubao-Seed-1.6 / Qwen3) and the three reasoning engines (HumanOmniV2 / Qwen2.5-Omni / Qwen2.5-VL), with other hyperparameters and inference configurations kept consistent to ensure comparability. Through these various combinations, we verify the magnitude of performance improvement when all three modules are integrated and assess the contribution of each module to overall performance.
Table 1. Summary of Models, Roles, and Scales (“a/b” denotes total / activated parameters, for Mixture-of-Experts models)
Model Role Parameter Scale
HumanOmniV2[7] Reasoning Engine 7B
Qwen2.5-Omni Reasoning Engine 7B
Qwen2.5-VL Reasoning Engine 7B
GPT-4o Strategy Generator/Strategy Selector Large Closed-Source Model
GLM-4.5 Strategy Generator/Strategy Selector 355B/32B
Doubao-Seed-1.6 Strategy Generator/Strategy Selector Large Closed-Source Model
Qwen3 Strategy Generator/Strategy Selector 235B/22B
Qwen2.5-VL-32B Intent Perceiver 32B
GLM-4.5V Intent Perceiver 106B/12B
Table2. Experimental Configurations: Three-Module Settings and Descriptions
Experiment ID Intent Perceiver Strategy Generation Strategy Selection Description
CG_Qwen_vl Qwen2.5-VL-32B 3 On All three modules enabled; use Qwen as Intent model
CG_GLM_vl GLM-4.5V 3 On All three modules enabled; use GLM as Intent model
Abl_NoIntent No 3 On Remove the Intent module
Abl_SingleStrategy Qwen2.5-VL-32B 1 No Change strategy generation to single strategy
BaseLine No No No No front-end pipeline
3.2 Main Results
Tables 3–5 report detailed results on the IntentBench, WorldSense, and Daily-Omni datasets, respectively. The “three-module” scheme consistently outperforms its respective baselines across all combinations of the four Pipelines and the three reasoning engines. The best results are sometimes achieved by CG_Qwen_vl (using Qwen2.5-VL-32B as the intent perceiver) and sometimes by the CG_GLM_vl (using GLM-4.5V as the intent perceiver), but both surpass “no intent” (Abl_NoIntent) and “single strategy” (Abl_SingleStrategy). In addition, both Abl_NoIntent and Abl_SingleStrategy outperform the baseline, confirming the effectiveness of each module.
Table 3. IntentBench dataset: accuracy (%) across combinations of reasoning models and Pipeline module models
Pipeline Reasoning Model (Baseline) CG_Qwen _vl CG_GLM _vl Abl_ NoIntent Abl_ Single Strategy
GPT-4o HumanOmniV2 (69.33) [7] 70.86 70.47 70.45 70.09
GLM-4.5 71.07 70.87 70.51 70.27
Doubao-Seed-1.6 70.92 70.72 70.06 69.74
Qwen3 71.18 70.9 70.82 69.96
GPT-4o Qwen2.5-Omni (64.2) [7] 65.95 66.07 65.82 64.99
GLM-4.5 65.51 65.86 65.31 65.31
Doubao-Seed-1.6 65.67 65.6 65.45 65.3
Qwen3 65.67 65.83 65.45 65.46
GPT-4o Qwen2.5-VL (61.68) 62.72 62.75 62.64 62.14
GLM-4.5 63.12 63.25 62.69 62.4
Doubao-Seed-1.6 63.2 63.02 62.9 62.09
Qwen3 63.81 63.83 63.78 63.39
Table 4. WorldSense dataset: accuracy (%) across combinations of reasoning models and Pipeline module models
Pipeline Reasoning Model (Baseline) CG_Qwen _vl CG_GLM _vl Abl_ NoIntent Abl_ Single Strategy
GPT-4o HumanOmniV2 (47.1) [7] 48.8 48.55 47.79 48.14
GLM-4.5 48.17 48.7 48.01 47.89
Doubao-Seed-1.6 48.23 48.2 48.14 47.64
Qwen3 48.36 48.36 48.3 47.92
GPT-4o Qwen2.5-Omni (45.4) [7] 47.13 47.67 47.01 46.75
GLM-4.5 47.57 47.38 47.23 46.31
Doubao-Seed-1.6 47.45 47.04 46.94 46.69
Qwen3 47.86 47.79 47.7 46.22
GPT-4o Qwen2.5-VL (37.39) 43.1 43.1 42.97 41.87
GLM-4.5 43.41 42.88 42.4 41.93
Doubao-Seed-1.6 42.21 42.12 41.93 41.33
Qwen3 43.06 43.25 42.91 41.83
Table 5. Daily-Omni dataset: accuracy (%) across combinations of reasoning models and Pipeline module models
Pipeline Reasoning Model (Baseline) CG_Qwen _vl CG_GLM _vl Abl_ NoIntent Abl_ Single Strategy
GPT-4o HumanOmniV2 (58.47) [7] 60.23 59.9 58.56 59.31
GLM-4.5 62.74 61.24 59.31 60.74
Doubao-Seed-1.6 62.66 61.07 59.9 60.9
Qwen3 62.49 61.32 61.24 60.15
GPT-4o Qwen2.5-Omni (47.45) [7] 55.56 55.64 55.47 50.38
GLM-4.5 54.05 54.22 53.38 50.88
Doubao-Seed-1.6 54.89 52.38 51.88 50.54
Qwen3 56.9 56.96 56.81 51.88
GPT-4o Qwen2.5-VL (47.28) 49.71 49.96 49.62 49.12
GLM-4.5 51.55 51.71 51.38 50.96
Doubao-Seed-1.6 50.79 50.63 50.46 50.35
Qwen3 51.63 51.55 51.46 50.54
From Table 3, on the IntentBench dataset: for HumanOmniV2, the baseline accuracy is 69.33%, and with our method it reaches a maximum of 71.18% (+1.85 pp; CG_Qwen_vl: Qwen2.5-VL-32B intent perceiver + 3 strategies + selector; Pipeline = Qwen3). For Qwen2.5-Omni, the baseline is 64.20%, with a maximum of 66.07% (+1.87 pp; CG_GLM_vl: GLM-4.5V intent perceiver + 3 strategies + selector; Pipeline = GPT-4o). For Qwen2.5-VL, the baseline is 61.68%, with a maximum of 63.83% (+2.15 pp; CG_GLM_vl; Pipeline = Qwen3).
On the WorldSense dataset (Table 4): HumanOmniV2 baseline 47.10%, maximum 48.80% (+1.70 pp; CG_Qwen_vl; Pipeline = GPT-4o); Qwen2.5-Omni baseline 45.40%, maximum 47.86% (+2.46 pp; CG_Qwen_vl; Pipeline = Qwen3); Qwen2.5-VL baseline 37.39%, maximum 43.41% (+6.02 pp; CG_Qwen_vl; Pipeline = GLM-4.5).
Similarly, on the Daily-Omni dataset (Table 5): HumanOmniV2 baseline 58.47%, maximum 62.74% (+4.27 pp; CG_Qwen_vl; Pipeline = GLM-4.5); Qwen2.5-Omni baseline 47.45%, maximum 56.96% (+9.51 pp; CG_GLM_vl; Pipeline = Qwen3); Qwen2.5-VL baseline 47.28%, maximum 51.71% (+4.43 pp; CG_GLM_vl; Pipeline = GLM-4.5).
In summary, the complete “three-module” intent-sketch reasoning scheme achieves stable improvements across all experiments (with a maximum gain of approximately 9.51%). Notably, these gains do not depend on a specific Pipeline model: even when the strategy generation and selection modules are replaced with smaller open-source models, large-model reasoning performance still exceeds the baseline without strategy prompts. This indicates that our method offers strong portability and plug-and-play characteristics, delivering consistent performance gains by embedding a unified strategy-prompting process regardless of the underlying model or platform interface.
3.3 Ablation Study and Analyses
To further understand the mechanisms of the three modules, we conduct ablations on Intent Perceiver, Strategy Generator, Strategy Selector. As shown in Tables 3–5, removing any single module leads to a drop in accuracy: taking Daily-Omni as an example, the complete HumanOmniV2 × GLM-4.5 configuration achieves 62.74%; removing the Intent Perceiver module (Abl_NoIntent) reduces it to 59.31% (−3.43 pp), while switching to a single strategy (Abl_SingleStrategy) yields 60.74% (−2.00 pp). Similar trends hold for other combinations: on WorldSense with Qwen2.5-VL × GLM-4.5, the complete configuration reaches 43.41%, while “no Intent Perceiver” gives 42.40% (−1.01 pp) and “single strategy” yields 41.93% (−1.48 pp); on IntentBench with HumanOmniV2 × Qwen3, the complete setting attains 71.18%, “no Intent Perceiver” 70.82% (−0.36 pp), and “single strategy” 69.96% (−1.22 pp). Without the intent provided by the Intent Perceiver, the model may miss crucial information, typically resulting in a modest performance drop; and when bypassing Strategy Selector to use a single strategy directly, the reasoning path may be suboptimal, leading to a more pronounced decline in accuracy. Notably, under some strong Pipeline configurations, the contribution of the Intent Perceiver module is relatively small (e.g., IntentBench with Qwen2.5-Omni × GPT-4o: complete 65.95% vs. no intent 65.82%, +0.13 pp), possibly because such powerful generative models already possess certain scene-understanding capabilities, partially offsetting the benefits of additional intent. Overall, “multi-strategy + selector” contributes the primary gains, while Intent Perceiver provides more substantial additional improvements in scenarios with stronger video/audio dependence (e.g., Daily-Omni). These ablation results indicate that the three-module component we propose performs distinct roles and that their synergy is necessary to maximally enhance reasoning performance: the Intent Perceiver supplies a clear scene context, the Strategy Generator broadens the solution-space of reasoning trajectories, and the Strategy Selector ensures the choice of the optimal approach, thereby making the final answers more accurate and reliable.
Our method does not depend on a specific large model or inference framework. As seen from the vertical comparisons in Tables 3–5, it consistently yields positive improvements across different reasoning engines (HumanOmniV2, Qwen2.5-Omni, Qwen2.5-VL) and different Pipelines (GPT-4o / GLM-4.5 / Doubao-Seed-1.6 / Qwen3). Moreover, for base models without any post-training, the relative gains from strategy prompting are more pronounced: for example, on WorldSense, Qwen2.5-VL improves from 37.39% to 43.41% (+6.02 pp; CG_Qwen_vl; Pipeline = GLM-4.5), and on HumanOmniV2 from 47.10% to 48.80% (+1.70 pp; CG_Qwen_vl; Pipeline = GPT-4o). This indicates that base models lacking post-training do not necessarily possess robust intent extraction and sound reasoning capabilities, and therefore benefit more from additional intent and strategy information. From the perspective of the Pipeline modules, the models we use span different “platforms,” including closed-source APIs and open-source models. The results show that regardless of which Pipeline model is adopted, the three-module strategy prompting has a positive effect on each reasoning engine: for example, on Daily-Omni, HumanOmniV2 improves from 58.47% to 62.74% (+4.27 pp; CG_Qwen_vl; Pipeline = GLM-4.5); even when replacing the Pipeline from GLM-4.5 (355B/32B) with the much smaller Qwen3 (235B/22B), the improvement for HumanOmniV2 remains close to four percentage points (+4.02 pp). This demonstrates that our method has strong platform-agnostic adaptability and generality: as a reasoning enhancement independent of underlying models, strategy prompting can be seamlessly embedded into different large-model inference engines to deliver cross-model performance gains.
Taken together, these findings show that the intent recognition, multi-strategy generation, and selection modules each contribute gains; the improvements do not rely on any particular reasoning engine or Pipeline; they are compatible with both closed- and open-source models of varying sizes; and the component operates in a plug-and-play manner, maintaining stable improvements across open-/closed-source settings and model scales, and under different engine/front-end substitutions—demonstrating cross-model, cross-platform, and cross-scenario generality and robustness.
CONCLUSION
This paper proposes a “human-like cognitive intent sketch” component that explicitly models the “understand–plan–select” reasoning process via a three-module pipeline (Intent Perceiver, Strategy Generator, Strategy Selector). As a plug-and-play, context-injected strategy, it improves the controllability and reliability of multimodal reasoning under a zero-shot setting without parameter fine-tuning, and offers strong portability and practicality across modalities, reasoning engines, and platforms. From an information-theoretic perspective, “multi-strategy generation + strategy selection” produces complementary candidates conditioned on the intent representation and selects the best by a minimum conditional-entropy/Bayesian-risk criterion, thereby effectively contracting the posterior over answers and reducing decision uncertainty and empirical error; this mechanism explains the stable gains and stronger interpretability achieved without any training overhead.
Experimental results show that on the IntentBench, WorldSense, and Daily-Omni benchmarks, the complete “three-module” scheme outperforms its respective baselines across all combinations of Pipelines and reasoning engines, with a maximum improvement of about +9.51 percentage points. Ablation analyses further indicate that the three modules are complementary and synergistic: “multi-strategy + selector” constitutes the primary source of gains, while explicit intent contributes especially in tasks with stronger video/audio dependence. Meanwhile, the method yields stable, cross-platform performance improvements for various reasoning engines and diverse Pipeline models, reflecting its plug-and-play operation and strong transferability. In sum, the “intent sketch” reasoning component provides a lightweight yet effective paradigm for enhancing the alignment, robustness, and interpretability of complex cross-modal reasoning.